OpenStudy (anonymous):

The minimum and maximum temperature for a day in Cupcake Town can be modeled by the equation below: 3|x − 8| + 15 = 18 What are the minimum and maximum temperatures for this day? A)x = 1, x = 17 B)x = −7, x = 9 C)x = 7 x = 9 D)No solution FAN/MEDALS

OpenStudy (anonymous):

A)x = 1, x = 17 B)x = −7, x = 9 C)x = 7 x = 9 D)No solution

pooja195 (pooja195):

Add -15 to both sides. 3(|x−8|)+15+−15=18+−15 3(|x−8|)=3 divide both sides by 3 3(|x−8|)/3=3/3 |x−8|=1 now set the solution to a psotivie and a negative value x−8=1 x−8=-1 solve both equations for x